OpenStudy (anonymous):

What is the 25th term in the following arithmetic sequence? -7, -2, 3, 8, ...

OpenStudy (ghazi):

T25= a+(n-1)d , where a is first term, d is common difference and n is nth term...you'll get T25= -7+(25-1)*5=.......

OpenStudy (ghazi):

what you got?

OpenStudy (anonymous):

T25=113

OpenStudy (anonymous):

thank you!